# Gradient-flow adaptive importance sampling for Bayesian leave one out cross-validation with application to sigmoidal classification models

### v2

@@ 4 Examples
The Hessian matrix of $\mu$, while non-zero, is sparse because all of the following identities hold:
$\nabla_{\mathbf{b}_{k}}\nabla_{\mathbf{b}_{j}}\mu$ $=\mathbf{0}\qquad\forall j,k$ (38)
$\nabla_{\mathbf{W}_{k}}\nabla_{\mathbf{W}_{k}}\mu$ $=\mathbf{0}\qquad\forall k$ (39)
$\nabla_{\mathbf{W}_{k}}\nabla_{\mathbf{b}_{j}}\mu$ $=\mathbf{0}\qquad\forall j\geq k.$ (40)
For this reason, the Jacobian determinant approximation of Eq. 27 can ignore the model Hessian entirely. However, in the case of one hidden layer we exploit the Hessian’s structure to provide explicit exact expressions for $\mathcal{J}_{(\cdot)}.$
@@
$\mathbf{v}_{k}^{\pm}=\left(\begin{matrix}\tilde{\mathbf{u}}_{k}/\sqrt{2|\mathbf{u}_{k}|^{2}}&\pm\mathbf{e}_{k}/\sqrt{2}&0\end{matrix}\right)^{\intercal},$ (43)
where
$\tilde{\mathbf{u}}_{k}=\big{(}\overbrace{0\quad\ldots\quad 0}^{(k-1)p\textrm{ zeros}}\quad\mathbf{u}_{k}^{\intercal}\quad\overbrace{0\quad\ldots\quad 0}^{(d-k)p\textrm{ zeros}}\big{)}^{\intercal},$ (44)
@@ 5 Experiments
PDF versions of Jupyter notebooks for producing the results in the text are included in the Supplemental Results. As baselines for comparison, we also evaluated generalizations of the moment matching (MM1/MM2) affine transformation method of Paananen et al. [2021] where we used a pre-factor of $\rho$ to control the magnitude of the change in the transformation (details in  S.3), and the log-likelihood (LL) gradient descent method of Elvira et al. [2022] (derivations of this transformation for sigmoidal models are available in  S.1).
Dataset and model: For demonstration, we used a public domain ovarian cancer micro-array dataset Hernández-Lobato et al. [2010], Schummer et al. [1999]. This dataset consists of $n=54$ observations of $p=1056+1$ predictors. As an example of a $p\gg n$ problem, model-agnostic $1/\ell$ importance sampling is insufficient for computing LOO expectations. Notably, Paananen et al. [2021] used this dataset to test their moment-matching adaptive importance sampler where they successfully decreased the number of observations where $\hat{k}>0.7$ from approximately $35$ to approximately $20$ using $s=1000$ posterior samples. We reproduced their logistic regression model, using the same regularized-horseshoe [Piironen and Vehtari, 2017b, c, Carvalho et al., 2009] prior, and the same statistical inference scheme within Stan, which we interfaced to Python using the package cmdstanpy. We ran four parallel Markov Chains, with twelve thousand burn-in iterations, retaining 2000 samples per chain (more details available in  S.4.1). We then evaluated the transformation methods on resamplings of the retained MCMC samples.
Adaptation:
KL Var LL MM1 MM2 #u̇nsuccessful
$31.5\pm 2.8$
$16.2\pm 3.0$
$18.2\pm 2.6$
$12.41\pm 2.6$
$20.4\pm 3.1$
$2.9\pm 1.6$
$3.1\pm 3.1$
$0.8\pm 1.0$
$0.2\pm 0.1$
Table 1: Counts of unsuccessful adaptations (mean $\pm$ standard deviation) when using at least one of the given combination of transformations across the step sizes $\rho\in\{4^{-r}:r\in\{0,1,\ldots,10\}\}$, as seen in one hundred simulations of parameter sample size $s=1000.$ Lower is better.
@@
We scanned different values of $\rho=4^{-r}$, for $r\in\{0,1,2,\ldots,10\}$, evaluating both transformations (KL/Var), and the comparison methods (LL/MM1/MM2) for a given value of $\rho.$ We performed this procedure $100$ times, using samples of size $s=1000.$ Recall that adaptation is successful if any of the considered transformations can reduce $\hat{k}$ to below $0.7$.
Table 1 presents statistics (mean $\pm$ standard deviation) for the number of observations where adaptation fails when using the given combination of methods. For a representative instance of the simulation procedure, Fig. 2 depicts the computed values of $\hat{k}$ obtained, organized by the index of each relevant observation within the dataset. The large black dot is the original $\hat{k}$ value and the smaller colored dots are the post-transformation KL and Var values for varying values of $\rho$. Fig. 3 shows the corresponding LOO- receiver operator characteristic (ROC) and precision recall curve (PRC) obtained by using the transformation that resulted in the best $\hat{k}$ value for each observation, predicting the LOO estimate of predictive probability, and feeding those probabilities into the relevant formulae for computing ROC and PRC. We provide these curves for both MCMC and ADVI -inferred variants of the model.
Figure 3: LOO ROC curves for ovarian cancer classification model contrasting the model fitted using MCMC and the model fitted using mean-field ADVI.
@@ 6 Discussion
Our methodology is based on taking a single step according to the gradient flow corresponding to the minimization of a given objective. We introduce two such transformations: KL divergence descent and variance descent. We presented explicit formulae for these transformations for logistic regression and ReLU-activated artificial neural networks with one hidden layer – the latter by computing the exact spectral decomposition of its Hessian matrix. An analysis of the same dataset using a Bayesian ReLU-net is available in S.4.2. We described how one can easily approximate the Jacobian of the transformations for more-complicated models, including for ReLU neural networks of any size. The adaptive importance sampler is ultimately used to estimate the expected LOO prediction for each given datapoint – quantities that can be used to compute downstream model generalization metrics such as ROC/PRC curves and the area under these curves.
@@
Contrasting and synergizing methods: Examining Tabel 1, taken individually, the KL and Var gradient flow-based transformations perform comparably to the original MM transformations evaluated in Paananen et al. [2021]. However, the generalized MM1/MM2 transformations have by-far the best performance in shifting $\hat{k}$. Yet these two MM methods, used either alone or together, usually do not completely get the job done. Each unsuccessful adaptation means that the model must be refit one additional time at high computational cost. Fortunately, using all the evaluated methods in-unison resulted in successful adaptation for all data points most of the time. The general strategy is then to loop through observations and try successive transformations for each observation until adaptation is successful.
@@
Limitations: The main tradeoff of this method versus the model-agnostic PSIS-LOO method is that this method is model-dependent. In order to use this methodology for a given model, one needs to be able to evaluate gradients of the model with respect to parameters – and also the gradients of the corresponding prior distribution. Both the KL descent and variance descent transformations require computing the the posterior density – when a variational approximation of the posterior is not available or trustworthy this computation is costly for large datasets.
@@
Extensions: In this manuscript, we focused on classification problems but the methodology for adapting the importance sampler is much broader. In the Supplemental Materials one may find more-general formulae for the KL and variance descending transformations. In medical and industrial contexts, one is often interested in whether an individual or unit will experience an outcome within a certain time interval. For instance, policymakers are interested in hospital readmission within 30 days post discharge [Xia et al., 2023, Chang et al., 2023] because these readmissions are possibly preventable. In these problems, one may apply survival modeling to characterize the lifetime distribution, and additionally evaluate a model according to its classification performance at a given cut-off time $T$. Our methodology can easily be used for assessing such models.
@@
Another extension to this methodology is to take more steps along the gradient flow for a given objective. It may be feasible to learn such a transformation using neural or other expressive representations. Finally, we resolved the spectrum of the Hessian matrix for shallow ReLU models – the spectral decomposition for larger ReLU models may be useful for other analyses.

### v3

@@ 4.2 Bayesian (ReLU) Neural Networks
The Hessian matrix of $\mu$, while non-zero, is sparse because all of the following identities hold: $\nabla_{\mathbf{b}_{k}}\nabla_{\mathbf{b}_{j}}\mu=\mathbf{0}\ \ \forall j,k,\quad\nabla_{\mathbf{W}_{k}}\nabla_{\mathbf{W}_{k}}\mu=\mathbf{0}\ \ \forall k,\quad\nabla_{\mathbf{W}_{k}}\nabla_{\mathbf{b}_{j}}\mu=\mathbf{0}\ \forall j\geq k.$ For this reason, the Jacobian determinant approximation of Eq. 22 can ignore the model Hessian entirely. However, in the case of one hidden layer we exploit the Hessian’s structure to provide explicit exact expressions for $\mathcal{J}_{(\cdot)}.$
@@
$\mathbf{v}_{k}^{\pm}=\left(\begin{matrix}\tilde{\mathbf{u}}_{k}/\sqrt{2|\mathbf{u}_{k}|^{2}}&\pm\mathbf{e}_{k}/\sqrt{2}&0\end{matrix}\right)^{\intercal},\quad\textrm{where}\quad\tilde{\mathbf{u}}_{k}=\big{(}\overbrace{0\quad\ldots\quad 0}^{(k-1)p\textrm{ zeros}}\quad\mathbf{u}_{k}^{\intercal}\quad\overbrace{0\quad\ldots\quad 0}^{(d-k)p\textrm{ zeros}}\big{)}^{\intercal},$ (32)
@@ 5 Experiments
Jupyter notebooks for producing the results in the text are included in the Supplemental Results. As baselines for comparison, we evaluated the original MM1/MM2 affine transformation methods of Paananen et al. (2021), and the log-likelihood (LL) gradient descent method of Elvira et al. (2022) (derivations of this transformation for sigmoidal models are available in  S.1). Note that in Paananen et al. (2021) they used a split-sampling scheme noting that all adaptations failed if it were omitted. In order to provide the most-direct comparison between the different transformations, we incorporated the MM1/MM2 transformations in the absence of split sampling.
5.1 Dataset and model
For demonstration, we used a public domain ovarian cancer micro-array dataset Hernández-Lobato et al. (2010); Schummer et al. (1999), consisting of $n=54$ observations of $p=1056+1$ predictors. As an example of a $p\gg n$ problem, model-agnostic $1/\ell$ importance sampling is insufficient for computing LOO expectations. Paananen et al. (2021) used this dataset to test their moment-matching adaptive importance sampler (in conjunction with split sampling) where they successfully decreased the number of observations where $\hat{k}>0.7$ from approximately $35$ to approximately $20$ using $s=1000$ posterior samples. We reproduced their logistic regression model, using the same regularized-horseshoe (Piironen & Vehtari, 2017b; c; Carvalho et al., 2009) prior, and the same statistical inference scheme within Stan, which we interfaced to Python using the package cmdstanpy. We ran four parallel Markov Chains, with twelve thousand burn-in iterations, retaining 2000 samples per chain (more details available in  S.3.1). We then evaluated the transformation methods on resamplings of the retained MCMC samples.
5.2 Adaptation
PMM1 PMM2 KL Var LL MM1 MM2 LR ovarian ReLUnet ovarian
- - - - - - - $34.9\pm 2.8$ $16.8\pm 3.7$
- - - - - - $5.3\pm 1.7$ $0.8\pm 0.9$
- - - - - - $5.3\pm 2.1$ $0.1\pm 0.3$
- - - - - - $17.9\pm 2.8$ $12.2\pm 2.3$
- - - - - - $18.4\pm 1.8$ $16\pm 3.4$
- - - - - - $22.2\pm 3.2$ $12.0\pm 3.0$
- - - - - - $34.8\pm 2.8$ $9.8\pm 2.9$
- - - - - - $34.9\pm 2.8$ $16.6\pm 3.7$
- - - $0.4\pm 0.5$ $0.0\pm 0$ 0.0
$0.3\pm 0.4$ $0.0\pm 0$ 0.0
This work Comparisons
Table 1: Counts of unsuccessful adaptations (mean $\pm$ standard deviation) when using at least one of the given combination of transformations across the step sizes $\bar{h}\in\{4^{-r}:r\in\{0,1,\ldots,10\}\}$, as seen in one hundred simulations of parameter sample size $s=1000.$ Lower is better.
Figure 2: Scatterplot of estimated Pareto tail shape diagnostic $\hat{k}$ versus observation, for transformed ovarian cancer logistic regression model parameters, for observations where the un-transformed samples have tail shape diagnostic $\hat{k}>0.7$ (black dots
@@ 5.2 Adaptation
Values of minimum $\hat{k}$ for each transformation plotted: green for KL, blue for Var, red for PMM1, orange for PMM2, purple for MM1, tan for MM2, and brown for LL – the minimum observed value for each transformation labeled. Adaptation for an observation is successful if $\hat{k}<0.7$ for any transformation. If the minimum value for a given transformation and observation falls outside of this displayed range then the corresponding point is omitted from this plot.
We scanned different values of $\bar{h}=4^{-r}$, for $r\in\{0,1,2,\ldots,10\}$, evaluating all transformations (KL/Var/PMM1/PMM2), and the comparison methods (LL/MM1/MM2) for a given value of $\bar{h}.$ We performed this procedure $100$ times, using samples of size $s=1000.$ Recall that adaptation is successful if any of the considered transformations can reduce $\hat{k}$ to below $0.7$.
Table 1 presents statistics (mean $\pm$ standard deviation) for the number of observations where adaptation fails when using the given combination of methods. When using all methods in unison, one is generally able to successfully prevent the need to refit either the logistic regression or the neural network models for the task of obtaining LOO statistics. In particular the PMM1/PMM2 methods were highly effective for the RELUnet model.
For a representative instance of the simulation procedure in the context of logistic regression, Fig. 2 depicts the minimum value of $\hat{k}$ obtained for each transformation, organized by the index of each relevant observation within the dataset. For each observation, the symbol
[…]
marks the pre-transformation value of $\hat{k}$. Generally, as seen in Fig. 2, PMM1/PMM2 had the most success. Hwoever, there are particular instances such as observations 23 and 46 where PMM1/PMM2 fail and the gradient flow transformations succeed in adaption. There are also many cases where multiple transformations each successfully adapt the posterior.
Fig. 3 shows the corresponding LOO- receiver operator characteristic (ROC) and precision recall curve (PRC) obtained by using the transformation that resulted in the best $\hat{k}$ value for each observation, predicting the LOO estimate of predictive probability, and feeding those probabilities into the relevant formulae for computing ROC and PRC. We contrast these curves for both MCMC and mean field ADVI -inferred variants of the model. The models inferred using MCMC have better generalization performance than their corresponding mean field approximations, which is not surprising due to the expected high degree of multicolinearity in this $p\gg n$ problem. This finding also held for the neural network model where it is notable that the MCMC-fitted neural network does not suffer as much from overfitting as does the mean field ADVI fitted neural network.
(a) LOO ROC curves for logistic regression ovarian cancer classification models contrasting the model fitted using MCMC and the model fitted using mean-field ADVI.
(b) LOO ROC curves for neural network ovarian cancer classification models contrasting the model fitted using MCMC and the model fitted using mean-field ADVI.
Figure 3: LOO ROC curves for ovarian cancer classification models fitted using MCMC and ADVI.
@@ 6 Discussion
Our methodology is based on taking samples from the posterior and transforming them by adding a small step either towards the target expectation or according to the gradient flow corresponding to the minimization of a given objective. We introduce four such transformations: our PMM1/PMM2 generalizations of the MM1/MM2 transformations in Paananen et al. (2021) and the KL/Var gradient flow step transformations. We presented explicit formulae for these transformations for logistic regression and ReLU-activated artificial neural networks. We described how one can easily approximate the Jacobian of the transformations for more-complicated models, including for ReLU neural networks of any size. The adaptive importance sampler is ultimately used to estimate the expected LOO prediction for each given datapoint – quantities that can be used to compute downstream model generalization metrics such as ROC/PRC curves and the area under these curves. Doing so without the need to refit a model saves considerable compute time and energy resources.
6.1 Contrasting and synergizing methods
@@
Examining Table 1, taken individually, the KL and Var gradient flow-based transformations perform comparably to the original MM1/MM2 transformations (in unison with split sampling) evaluated in Paananen et al. (2021). Of note, as in Paananen et al. (2021), MM1/MM2 used in the absence of splitting was unable to successfully adapt any observation in our evaluations. However, the generalized PMM1/PMM2 transformations have by-far the best performance in shifting $\hat{k}$. Yet these two PMM methods, used either alone or together, usually do not completely get the job done. Each unsuccessful adaptation means that the model must be refit one additional time at high computational cost. Fortunately, using all the evaluated methods in-unison resulted in successful adaptation for all data points most of the time. The general strategy is then to loop through observations and try successive transformations for each observation until adaptation is successful.
6.2 Limitations
@@
The main tradeoff of this method versus the model-agnostic PSIS-LOO method is that this method is model-dependent. In order to use this methodology for a given model, one needs to be able to evaluate gradients of the model with respect to parameters – and also the gradients of the corresponding prior distribution. Both the KL descent and variance descent transformations require computing the the posterior density – when a variational approximation of the posterior is not available or trustworthy this computation is costly for large datasets.
6.3 Extensions
@@
In this manuscript, we focused on classification problems but the methodology for adapting the importance sampler is much broader. In the Supplemental Materials one may find more-general formulae for the KL and variance descending transformations. In medical and industrial contexts, one is often interested in whether an individual or unit will experience an outcome within a certain time interval. For instance, policymakers are interested in hospital readmission within 30 days post discharge (Xia et al., 2023; Chang et al., 2023) because these readmissions are possibly preventable. In these problems, one may apply survival modeling to characterize the lifetime distribution, and additionally evaluate a model according to its classification performance at a given cut-off time $T$. Our methodology can easily be used for assessing such models.
@@
Another extension to this methodology is to take more steps along the gradient flow for a given objective. It may be feasible to learn such a transformation using neural or other expressive representations.